# Training Language Models to Explain Their Own Computations

### v2

@@ Abstract
Can language models (LMs) learn to faithfully describe their internal computations? Are they better able to describe themselves than other models? We study the extent to which LMs’ privileged access to their own internals can be leveraged to produce new techniques for explaining their behavior. Using existing interpretability techniques as a source of ground truth, we fine-tune LMs to generate natural language descriptions of (1) the information encoded by LM features, (2) the causal structure of LMs’ internal activations, and (3) the influence of specific input tokens on LM outputs. When trained with only tens of thousands of example explanations, explainer models exhibit non-trivial generalization to new queries. This generalization appears partly attributable to explainer models’ privileged access to their own internals: using a model to explain its own computations generally works better than using a different model to explain its computations (even if the other model is significantly more capable). Our results suggest not only that LMs can learn to reliably explain their internal computations, but that such explanations offer a scalable complement to existing interpretability methods. [^1] Code and data at https://github.com/TransluceAI/introspective-interp.
@@ 1 Introduction
When language models (LMs) are prompted to explain decisions—to describe what features they relied on, or what intermediate computations they performed when generating an answer—their outputs are often plausible (korbak2025chainthoughtmonitorabilitynew). But even when plausible, these explanations may not have any relationship to the internal computation that gave rise to LMs’ decisions in the first place. Indeed, work using attribution, feature visualization, and mechanistic interpretability techniques (Simonyan2013DeepIC; 10.5555/3495724.3497168; templeton2024scaling) has shown that LMs’ verbalized explanations sometimes systematically fail to describe the factors that determine their decisions (turpin2023language; chen2025reasoningmodelsdontsay; barez2025cot).
Can we fix this problem, and train LMs so that their verbalized explanations faithfully describe their internal computations? Several recent papers (binder2025looking; song2025privilegedselfaccessmattersintrospection; plunkett2025selfinterpretabilityllmscomplexinternal) have studied the related question of whether models can describe features of their own output distributions, arguing that models might be able to leverage privileged access to their own internals to do so effectively. In this paper, we are interested in an even deeper form of privileged access: whether models can learn to describe not only their outputs, but their internal representations and mechanisms, and whether privileged self-access enables them to do so better than learned explanation-generating methods derived from other models. We state our main hypothesis as follows:
@@
    Descriptions of internal features, specifically the kinds of inputs that activate them (feature descriptions; Figure˜1A).
@@
    Descriptions of decision rules, specifically the important tokens for a decision (input ablation; Figure˜1C). We focus on a setting where recent work found that reasoning models failed to produce faithful chains-of-thought (chen2025reasoningmodelsdontsay), in order to explore whether training can improve faithfulness. This setting also bears similarities to binder2025looking due to measuring self-access to the model’s output distributions.
Our main experiments focus on feature descriptions, where we find evidence of privileged access—our target model is better explained by itself than by other explainer models, even when the other model is generally more capable (e.g. larger or instruction-tuned).
We dive deeper into this phenomenon and probe the generality of this hypothesis, finding that:
@@
    LMs can be fine-tuned to self-explain: As in past work (li2025do), we find that explanation capabilities are not always natively present in models, and must be acquired via fine-tuning. A small amount of fine-tuning enables LMs to faithfully describe their internal features, significantly outperforming zero-shot methods (kharlapenko2024self; chen2024selfie) and nearest-neighbor SAE features (huben2024sparse). Furthermore, a model’s ability to explain (its own or others’) features is correlated with the degree of similarity between the explainer and target models (§3.4).
@@
    Self-explaining is data-efficient: Self-explaining is approximately a hundred times more sample-efficient than a nearest neighbors baseline, achieving comparable results with only 0.8% of the training data (§4). This makes self-explaining a scalable approach to generating feature descriptions.
@@
    Privileged access extends to other tasks: we train models to explain the effects of activation patching (§5.1) and input ablation (§5.2). For the first task, we see the same pattern of privileged access as in feature descriptions. For the second, we see a slightly different pattern, as some models appear to be universally stronger explainers (e.g. Qwen outperforms Llama across the board); however, we still observe evidence of privileged access here (e.g. Qwen is better at explaining itself than explaining Llama, and vice versa).
Taken together, these results suggest that even when language models cannot faithfully self-explain as a result of ordinary training, they can learn to do so through an objective that enforces consistency between their external explanations and their internal procedures. This reframes interpretation as not only an external analysis problem, but as a capability that can be trained into LMs themeselves; by leveraging privileged access to internal computations, “introspective interpretability” techniques offer an avenue towards scalable understanding of model behavior.
@@ 2.1 Definitions and Task
Abstractly, a neural model $\mathcal{M}$ can be represented as a pipeline $x\to h\to y$ that maps inputs $x$ to hidden representations $h$ to outputs $y$. An explanation $E$ of the neural network may enable us to answer questions $q$ about what information in $x$ is represented in $h$, how that information influences $y$, and what pieces of the model are responsible for each step.
For example, we may wish to answer a question $q=\textit{What inputs activate direction $v$ at layer $l$?}$ If $v$ activates highly on tokens Tokyo, NYC, and Paris, then an appropriate explanation $E$ might be city names. We are interested in whether LMs can learn to generate them directly.
@@
Our experiments focus on explaining Transformer language models $\mathcal{M}$ (grattafiori2024llama3herdmodels; yang2025qwen3technicalreport), which consist of layers $\ell$ each containing attention and MLP blocks, and residual stream representations $h_{\ell}$ in between each layer. We also use Transformer language models as explainers $\mathcal{E}$. Throughout this paper, we use notation $x=(x_{1},\cdots x_{n})$ to refer to an input sequence with token positions $t\in[n]$. We use $h_{(\ell,t)}(x)$ to denote the residual stream activation of $\mathcal{M}$ at layer $\ell$ and token position $t$ for input $x$.
@@ 2.2 Feature Descriptions: What does a model-internal feature represent?
Figure 2: Training an explainer to predict feature descriptions. (A) We run a target model $\mathcal{M}$ across many input contexts. (B) We then observe vectors $v$ at layers $\ell$ and select an explanation $E$ that best describes the contexts in which $v$ is active, which we assess via a LM simulator. (C) Finally, we train an explainer model to respond with $E$ when given the question $q$ about what $v$ means at layer $\ell$, and find that it performs well on both held-out and OOD vectors $v$.
@@
For feature description, the question $q$ is what types of inputs activate direction $v$ in the residual stream at layer $l$?
@@
We make use of existing automated interpretability methods (hernandez2022natural; bills2023language; choi2024automatic), which generate natural-language descriptions $E$ by summarizing input tokens on which a given $v$ is highly active.
Formally, let $a_{v}(x,\ell,t)=\langle h_{\ell,t}(x),v\rangle$ denote the activation of $v$ in the residual stream representation of the token $x_{t}$. To obtain feature descriptions, we begin by prompting (or training) a “simulator” language model (LM simulator in Figure˜2) to map from candidate descriptions $E$ (e.g. Encodes city names in Figure˜2) and input sequences $x$ (e.g. we’ll go visit NYC) to predicted activations $\hat{a}(x,t,E)$ (e.g. outputting 0.1 for visit and 0.9 for NYC). See Figure˜2B. We train the simulator following choi2024automatic on FineWeb exemplars (penedo2024the) with Neuronpedia SAE labels.
Given the simulator, for each direction $v$, we search for the explanation $E$ that maximizes correlation between simulated and true activations over all input data:
$T_{\text{feat}}(\mathcal{M},v,\ell)=\operatorname*{arg\,max}_{E}\mathbb{E}_{x}\left[\texttt{corr}_{t}\left(a_{v}(x,\ell,t),\,\hat{a}(x,t,E)\right)\right]~.$ (2)
Intuitively, this procedure will assign each feature $v$ an explanation $E$ that describes the tokens on which $v$ is most active. The prompts we use to encode $q$ and $E$ are given in Appendix˜D.
@@
$\mathcal{E}$ receives a question $q$ that contains a feature $v$ and a layer $\ell$, and is trained to generate the explanation $E$ of when $v$ is highly active at layer $\ell$. We take $v$ from the target model $\mathcal{M}$ at its layer-$\ell$ residual stream, and pass it as a continuous token to the explainer, inserted at the embedding layer. This enables the explainer to accept arbitrary vectors vv. [^3] In early experiments, we also investigated explaining features $v$ from other components of the target model, but found that explainers struggled to reproduce these explanations. We hypothesize that we need to patch features into the same type of location (i.e. the embedding layer is part of the residual stream). See Figure˜2C.
For explainer models whose hidden dimension does not match the target model, we introduce a linear projection $\Pi_{\ell}\in\mathbb{R}^{d_{\mathcal{E}}\times d_{\mathcal{M}}}$ from the target model hidden size $d_{\mathcal{M}}$ to the explainer model hidden size $d_{\mathcal{E}}$, which is trained jointly with the rest of LM parameters. We learn a separate projection per layer $\ell$ of the target model. We train explainers and projections to optimize:
$\boxed{\mathcal{L}_{\text{feat}}=-\mathbb{E}_{v,\ell}\left[\log p_{\mathcal{E}}(E=T_{\text{feat}}(\mathcal{M},v,\ell)\mid\Pi_{\ell}\,v,\ell)\right]}$ (3)
@@
Because activations $h$ tend to be highly polysemantic (elhage2022superposition), interpretation is often done on decomposed representations, which are commonly either individual neurons (so $v$ is a one-hot vector; bau2020units) or Sparse Auto-Encoder (SAEs) features obtained from sparse dictionary learning approaches (with $v$ being a member of the learned dictionary; huben2024sparse; bricken2023monosemanticity). Our experiments train on SAE features, but test generalization to full activations and activation differences. See details in Section˜3.2.
@@ 2.3 Activation Patching: What internal components affect the prediction?
Figure 3: Training an explainer to predict activation patching outcomes. (A) We run a target model $\mathcal{M}$ on an input $x$ and obtain a prediction $\mathcal{M}(x)$. (B) We perform activation patching by running $\mathcal{M}$ on a counterfactual input $x^{\prime}$—in this example, patching in vector $v$ from token $x_{1}$ and layer $\ell_{1}$ of the counterfactual run into the original run—and then construct an explanation $E$ about how the resulting prediction changes. (C) Finally, we train an explainer model to answer $E$ when given a question $q$ about the patching procedures.
@@
For activation patching, we ask which hidden components are causally important for a model’s output (meng2022locating; zhang2024towards). Specifically, the question $q$ is on input $x$, (how) does changing the activation at layer $\ell$ and token $x_{t}$ affect the output? An example is shown in Figure˜3.
@@
We can answer questions of this form through activation patching and circuit tracing methods. Given an input $x$ (Paris is the capital of in Figure˜3), an activation patching experiment typically constructs an alternative input $x^{\prime}$ (Rome is the capital of). During the forward pass, we replace the activation $h_{\ell,t}(x)$ with the corresponding activation $h_{\ell,t}(x^{\prime})$:
@@
$T_{\text{patch}}(\mathcal{M},q)=T_{\text{patch}}(\mathcal{M},x,\ell,t)=d\left(\mathcal{M}(x),\;\mathcal{M}\left(x;\,h_{\ell,t}(x)\leftarrow h_{\ell,t}(x^{\prime})\right)\right),$ (5)
where $d(\cdot,\cdot)$ is some measure of the difference between the two predictions. This process is illustrated in Figure˜3B.
@@
Additionally, rather than patching activations one layer at a time, we modify representations at several layers simultaneously to ensure sufficient samples where the model’s output changed under intervention. We divide the target model’s layers into four blocks, and when performing activation patching, compute an aggregate representation $v=\texttt{avg}_{\ell_{k}}\left(h_{\ell_{k},t}(x^{\prime})\right)$ from the counterfactual input, then insert $v$ at all corresponding layers of the original input. [^4] We perform aggregation so that we only have to specify a single continuous token to the explainer. Finally, we optimize:
$\boxed{\mathcal{L}_{\text{patch}}=-\mathbb{E}_{x,x^{\prime},t,\ell_{1\cdots i}}\left[\log p_{\mathcal{E}}\left(E=T_{\text{patch}}(\mathcal{M},x,t,\ell_{1\cdots i},v)\mid x,t,\ell_{1\cdots i},v\right)\right]}$ (6)
The exact prompt and output templates are provided in Appendix˜D.
@@ 2.4 Input Ablation: What parts of the input matter?
Figure 4: Training an explainer to predict input ablation outcomes. We run a target model $\mathcal{M}$ on an input $x$ (A), then identify whether a part of the input $\tilde{x}\subset x$ was important to the output by rerunning $\mathcal{M}$ on $\x~x\backslash\tilde${x} (B). Finally, we train an explainer model to answer question $q$ about how the output changes when $\tilde{x}$ is removed from $x$ (C).
@@
For input ablation, we ask which input tokens most influence the model’s prediction (chang-etal-2025-jopa; vafa-etal-2021-rationales). Specifically, the question $q$ is Are tokens $\tilde{x}$ in $x$ important to the model’s prediction? For instance, in Figure˜4, the question $q$ asks whether the hint is important to the answer.
@@
As with activation patching, we can remove (subsets of) tokens in $x$ and identify which ones cause predictions $y$ to change. Let $\tilde{x}\subseteq x$ be a subset of tokens to remove. Input ablation measures how much the model output changes when we take an input $x$ (e.g. The first US president… in Figure˜4) and remove a subset of tokens $\tilde{x}$ (e.g. Hint: B):
@@
$\mathcal{E}$ receives a question $q$ that contains a multiple-choice reasoning question $c$ with a hint $\tilde{x}$, and is asked to predict (1) whether removing $\tilde{x}$ would change the model’s answer (the has-changed value), and (2) the new output without the hint (the content of the change). See Figure˜4C. Formally, the objective we optimize is:
$\boxed{\mathcal{L}_{\text{input}}=-\mathbb{E}_{x,\tilde{x}}\left[\log p_{\mathcal{E}}(E=T_{\text{input}}(\mathcal{M},x,\tilde{x})\mid x,\tilde{x})\right]}$ (8)
Full prompt and output templates are provided in Appendix˜D.

### v3

@@ Abstract
Can language models (LMs) learn to faithfully describe their internal computations? Are they better able to describe themselves than other models? We study the extent to which LMs’ privileged access to their own internals can be leveraged to produce new techniques for explaining their behavior. Using existing interpretability techniques as a source of ground truth, we fine-tune LMs to generate natural language descriptions of (1) the information encoded by LM features, (2) the causal structure of LMs’ internal activations, and (3) the influence of specific input tokens on LM outputs. When trained with only tens of thousands of example explanations, explainer models exhibit non-trivial generalization to new queries. This generalization appears partly attributable to explainer models’ privileged access to their own internals: fine-tuning a model to explain its own computations generally works better than fine-tuning a different model (even if the explainer model is significantly more capable than the target). Our results suggest not only that LMs can learn to reliably explain their internal computations, but that such explanations offer a scalable complement to existing interpretability methods. [^1] Code available at https://github.com/TransluceAI/introspective-interp.
Machine Learning, ICML
@@ 1 Introduction
When language models (LMs) are prompted to explain decisions, their outputs are often plausible (Korbak et al., 2025), but may not bear any relationship to the computation that first gave rise to LMs’ decisions. Indeed, work using attribution, feature visualization, and mechanistic interpretability techniques (Simonyan et al., 2013; Covert et al., 2020; Templeton et al., 2024) has shown that LMs’ verbalized explanations sometimes systematically fail to describe the factors that determine their decisions (Turpin et al., 2023; Chen et al., 2025; Barez et al., 2025).
Can we fix this problem by training LMs so that their verbalized explanations faithfully describe their internal computations? Several recent papers (Binder et al., 2025; Song et al., 2025b; Plunkett et al., 2025) have studied the related question of whether models can describe features of their own output distributions, arguing that models might be able to leverage privileged access to their own internals to do so effectively. In this paper, we are interested in an even deeper form of privileged access: whether models can learn to describe not only their outputs, but their internal representations and mechanisms, and whether privileged self-access enables them to do so better than learned explanation-generating methods derived from other models. We state our main hypothesis as follows:
@@
    Descriptions of internal features, specifically what inputs activate them (feature descriptions; Figure˜1A).
@@
    Descriptions of decision rules, specifically the important tokens for a decision (input ablation; Figure˜1C). This setting bears similarities to Binder et al. (2025) due to measuring self-access to the model’s output distributions.
Our experiments find evidence of privileged access—the target model is better explained by itself than by other explainer models, even when the other model is more capable (e.g. larger or instruction-tuned). Diving deeper, we find:
@@
    LMs can be fine-tuned to self-explain: As in past work (Li et al., 2025), we find that explanation capabilities are not always natively present in models, and must be acquired via fine-tuning. A small amount of fine-tuning enables LMs to faithfully describe their internal features, significantly outperforming zero-shot methods (Kharlapenko et al., 2024; Chen et al., 2024a) and nearest-neighbor SAE features (Huben et al., 2024). Furthermore, a model’s ability to explain (its own or others’) features is correlated with the degree of similarity between the explainer and target models (§3.4).
@@
    Self-explaining is data-efficient: Self-explaining is approximately a hundred times more sample-efficient than a nearest neighbors baseline, achieving comparable results with only 0.8% of the training data (§4).
@@
    Privileged access extends across tasks: we find patterns of privileged access in all three domains, including activation patching (§5.1) and input ablation (§5.2).
These results suggest that even when off-the-shelf LMs cannot faithfully self-interpret, they can learn to do so through an objective that enforces consistency between their output and their internal procedures. Our approach reframes interpretability as not only an external analysis problem, but as a capability that can be trained into LMs themeselves; by leveraging privileged access to internal computations, “introspective interpretability” techniques offer an avenue towards scalable understanding of LM behavior.
@@ 2.1 Definitions and Task
Abstractly, a neural model $\mathcal{M}$ can be represented as a pipeline $x\to h\to y$ that maps inputs $x$ to hidden representations $h$ to outputs $y$. An explanation $E$ of the neural network may enable us to answer questions $q$ about what information in $x$ is represented in $h$ and how that information influences $y$. For example, we may wish to answer a question $q=\textit{What inputs activate direction $v$ at layer $l$?}$ If $v$ activates on tokens Tokyo, NYC, and Paris, then an appropriate explanation $E$ might be city names. We are interested in whether LMs can learn to generate $E$ directly.
@@
Our experiments focus on explaining Transformer language models $\mathcal{M}$ (Grattafiori et al., 2024; Yang et al., 2025), which consist of layers $\ell$ each containing attention and MLP blocks, and residual stream representations $h_{\ell}$ in between each layer. We also use Transformer language models as explainers $\mathcal{E}$. Throughout this paper, we use notation $x=(x_{1},\cdots x_{n})$ to refer to an input sequence with token positions $t\in[n]$. We use $h_{(\ell,t)}(x)$ to denote the residual stream activation of $\mathcal{M}$ at layer $\ell$ and token position $t$ for input $x$.
@@ 2.2 Feature Descriptions: What does a model-internal feature represent?
Figure 2: Training an explainer to predict feature descriptions. (A) Run the target model $\mathcal{M}$ on many inputs. (B) For each vector $v$ at layer $\ell$, pick an explanation $E$ that best matches the context where $v$ is active, as judged by a LM simulator. (C) Train the explainer to output $E$ given the question $q$ about $v$ at layer $\ell$. The trained explainer generalizes well to both held-out and OOD $v$.
@@
For feature descriptions, our question $q$ takes form what types of inputs activate direction $v$ in the residual stream at layer $l$?
@@
We make use of automated interpretability methods (Hernandez et al., 2022; Bills et al., 2023), which generate natural-language descriptions $E$ of input tokens on which a given $v$ is highly active.
Formally, let $a_{v}(x,\ell,t)=\langle h_{\ell,t}(x),v\rangle$ denote the “activation” of $v$ in the residual stream representation of token $x_{t}$ (i.e. the alignment between $h_{\ell,t}$ and $v$). To obtain feature descriptions, we use a “simulator” language model to simulate how $v$ would activate on an input sequence $x$ if it had description $E$, outputting an “expected” activation pattern $\hat{a}(x,t,E)$. See Figure˜2B. We train the simulator following Choi et al. (2024) using Neuronpedia SAE labels and FineWeb exemplars (Penedo et al., 2024).
We then search for the explanation $E$ that maximizes correlation between simulated and true activations over inputs:
$T_{\text{feat}}(\mathcal{M},v,\ell)=\operatorname*{arg\,max}_{E}\mathbb{E}_{x}\left[\texttt{corr}_{t}\left(a_{v}(x,\ell,t),\,\hat{a}(x,t,E)\right)\right]$ (2)
The prompts to encode $q$ and $E$ are given in Appendix˜F.
@@
We take $v$ from the layer-$\ell$ residual stream of the target model and pass it as a continuous token to the explainer by inserting $v$ at the embedding layer of the explainer model. [^3] In early experiments, we also investigated explaining features $v$ from other components of the target model, but found that explainers struggled to reproduce these explanations. We hypothesize that we need to patch features into the same type of location (i.e. the embedding layer is part of the residual stream). See Figure˜2C.
For explainer models whose hidden dimension do not match the target model’s, we introduce a linear projection $\Pi_{\ell}\in\mathbb{R}^{d_{\mathcal{E}}\times d_{\mathcal{M}}}$ from the target model hidden size $d_{\mathcal{M}}$ to the explainer model hidden size $d_{\mathcal{E}}$, which is trained jointly with the rest of LM parameters. We learn a separate projection per layer $\ell$ of the target model. We train explainers and projections to optimize:
$\hskip-10.00002pt\boxed{\mathcal{L}_{\text{feat}}=-\mathbb{E}_{v,\ell}\left[\log p_{\mathcal{E}}(E=T_{\text{feat}}(\mathcal{M},v,\ell)\mid\Pi_{\ell}\,v,\ell)\right]}$ (3)
@@
Our experiments train on Sparse Auto-Encoder (SAEs) features obtained from sparse dictionary learning approaches Huben et al., 2024; Bricken et al., 2023), but test generalization to full activations and activation differences. See details in Section˜3.2.
@@ 2.3 Activation Patching: What internal components affect the prediction?
Figure 3: Training an explainer to predict activation patching outcomes. (A) Run target model $\mathcal{M}$ on input $x$ to obtain $\mathcal{M}(x)$. (B) Perform activation patching by running $\mathcal{M}$ on a counterfactual input $x^{\prime}$—in this example, patching in vector $v$ from token $x_{1}$ and layer $\ell_{1}$ of the counterfactual run into the original run—and then construct an explanation $E$ about how the resulting prediction changes. (C) Train an explainer model to answer $E$ when given a question $q$ about the patching procedures.
@@
We ask which hidden components are causally important for a model’s output (Meng et al., 2022; Zhang & Nanda, 2024). Specifically, the question $q$ is on input $x$, (how) does changing the activation at layer $\ell$ and token $x_{t}$ affect the output? An example is shown in Figure˜3.
@@
We can answer questions of this form through activation patching and circuit tracing methods. Given an input $x$ (Paris is the capital of in Figure˜3), an activation patching experiment typically constructs an alternative input $x^{\prime}$ (Rome is the capital of). We then run a forward pass of $\mathcal{M}$ on $x$, but replace the activation $h_{\ell,t}(x)$ with the corresponding activation $h_{\ell,t}(x^{\prime})$:
@@
$T_{\text{patch}}(\mathcal{M},q)$ $=T_{\text{patch}}(\mathcal{M},x,\ell,t)$ (5)
$=d\Big(\mathcal{M}(x),\;\mathcal{M}\big(x;\,h_{\ell,t}(x)\leftarrow h_{\ell,t}(x^{\prime})\big)\Big).$
where $d(\cdot,\cdot)$ is some measure of the difference between the two predictions. This process is illustrated in Figure˜3B. [^4] To ensure sufficient samples where the model’s output changes We divide the target model’s layers into four blocks, and perform activation patching over one block of layers at a time We aggregate the representations over the block $v=\texttt{avg}_{\ell_{k}}\left(h_{\ell_{k},t}(x^{\prime})\right)$ before inserting into all corresponding layers of the original input, allowing us to specify only a single continuous token to the explainer.
@@
Finally, we optimize:
$q\triangleq(x,t,\ell_{1\cdots i},v).$ (6)
$\boxed{\mathcal{L}_{\text{patch}}=-\mathbb{E}_{x,x^{\prime},q}\Big[\log p_{\mathcal{E}}\big(E=T_{\text{patch}}(\mathcal{M},q)\mid q\big)\Big]}$ (7)
The exact prompt and output templates are in Appendix˜F.
@@ 2.4 Input Ablation: What parts of the input matter?
Figure 4: Training an explainer to predict input ablation outcomes. (A) Run a target model $\mathcal{M}$ on an input $x$. (B) Identify whether input component $\tilde{x}\subset x$ affected its output by rerunning $\mathcal{M}$ on $\x~x\backslash\tilde${x}. (C) Train an explainer model to answer question $q$ about how the output changes when $\tilde{x}$ is removed from $x$.
@@
For input ablation, we ask which input tokens most influence the model’s prediction (Chang et al., 2025; Vafa et al., 2021), which we can identify by ablating subsets $\tilde{x}$ of the input $x$. Specifically, $q$ takes form if $\tilde{x}$ is removed, (how) does the answer to $x$ change? See Figure˜4.
@@
Input ablation measures how much the target model $\mathcal{M}$’s output changes when we take an input $x$ (e.g. The first US president… in Figure˜4) and remove a subset of tokens $\tilde{x}\subseteq x$ (e.g. Hint: B):
@@
$\mathcal{E}$ is asked to predict (1) whether removing $\tilde{x}$ would change the $\mathcal{M}$’s answer, and (2) the content of $\mathcal{M}$’s new output without $\tilde{x}$. See Figure˜4C. Formally, the objective we optimize is:
$\hskip-10.00002pt\boxed{\mathcal{L}_{\text{input}}=-\mathbb{E}_{x,\tilde{x}}\left[\log p_{\mathcal{E}}(E=T_{\text{input}}(\mathcal{M},x,\tilde{x})\mid x,\tilde{x})\right]}$ (9)
Full prompt and output templates are in Appendix˜F.